OpenStudy (anonymous):

amplitude: 2 Period: 3pi Phase Shift: -1/9. What is the equation?

OpenStudy (aum):

In y = Acos(Bx + C) OR y = Asin(Bx + C) Amplitude = A Period = (2*pi) / B Phase Shift = -C/B Plug in the numbers and find A,B and C. Substitute A, B and C either in the cosine function or the sine function above.